OpenStudy (millsemily):

Hiya! Can someone check my answer? The student council has decided to sell t-shirts for a fundraiser. The printing company charges a $75 set-up fee plus $3 per shirt. The student council plans to sell their t-shirts for $12 per shirt. Write a function p(t) to represent the student council's profits. A) p(t) = 12t B) p(t) = 3t + 75 C) p(t) = 9t + 75 <--- D) p(t) = 9t - 75

OpenStudy (anonymous):

where is your work?

OpenStudy (anonymous):

Yes I think you're right because all the rest of them looks wrong. I did that exact same problem in 8th grade

OpenStudy (millsemily):

I did the work in my notebook already. I need someone to check my answer :) Just tell me if it's right or wrong, and I'll re-do the probably. I don't need the answer :)

OpenStudy (millsemily):

*problem

OpenStudy (anonymous):

exactly @Jayla_McManus yes you are corecct @millsemily

OpenStudy (aaronandyson):

YOUR CORRECT :)

OpenStudy (anonymous):

yes.

OpenStudy (millsemily):

Thank you :)

OpenStudy (anonymous):

np

OpenStudy (aaronandyson):

np:)

OpenStudy (ganpat):

Total Profit = sp - cp let t t-shirts be sold.. therefore, sp = 12t and cp = (75+3t) p(t) = 12t - (75+3t)... and you can get the answer !

OpenStudy (ganpat):

I meant you are wrong :p and the question is right :D